OpenStudy (anonymous):

Square root of X (X+1) Derivatives help

OpenStudy (anonymous):

Is it \[\sqrt{x}(x+1)\] or \[\sqrt{x(x+1)}\] ?

OpenStudy (anonymous):

√x (x+1)

OpenStudy (anonymous):

Ok, so \[x^{1/2}(x+1)\]

OpenStudy (anonymous):

I'd probably multiply it out, then take the derivative.

OpenStudy (anonymous):

If it is all under the radical then you'll need to use the chain rule instead.

OpenStudy (anonymous):

Yes you are right on both cases.

OpenStudy (anonymous):

My bad lol. I thought you were the one needing help polpak haha. Sorry.

OpenStudy (anonymous):

\[\frac{d}{dx}[ x^{3/2} + x^{1/2}] = \frac{3x^{1/2}}{2} + \frac{1}{2x^{1/2}}\] Then simplify.

OpenStudy (anonymous):

how did u get x^3/2+X^1/2

OpenStudy (anonymous):

\[x^{1/2}(x+1) = x^{1/2}*x^1 + x^{1/2}*1 = x^{\frac{1}{2}+1} +x^{1/2}\] \[ = x^{3/2} +x^{1/2}\]